# Leak Proof PDBBind: A Reorganized Dataset of Protein-Ligand Complexes for More Generalizable Binding Affinity Prediction

### v2

@@ 2.4 Preparation of two target-specific datasets
Finally, we have prepared two additional datasets of protein-ligand complexes with the same protein and different ligands so that we could evaluate the ranking accuracies of the SFs before and after retraining using LP-PDBBind. The first dataset is based on the SARS-CoV-2 main protease (Mpro) for which a wide variety of potential Mpro inhibitors have been developed.41 We have manually extracted published co-crystal structures of Mpro with a number of non-covalent inhibitors53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, and prepared a dataset containing 40 structures and corresponding experimental binding affinity measurements. The second dataset involves the epidermal growth factor receptor (EGFR), which is a receptor tyrosine kinase related to multiple cancers including lung cancer, pancreatic cancer and breast cancer.42 Similarly, we have selected 23 representative non-covalent protein-ligand complex structures of EGFR with binding affinities taken from BindingDB.66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80 The two dataset were selected as two representative scenarios of using the SFs on a similar or different protein-ligand complex than what was included in the training dataset.
@@ 2.5 Retraining Scoring Functions with LP-PDBBind
The LP-PDBBind CL1 cleaned data for non-covalent binders of PDBBind were used to retrain AutoDock Vina3, IGN15, the 2010 RF-Score14 and DeepDTA18. These models cover a wide range of different approaches for scoring, and are representative of different dimensions of the SF space of models. AutoDock vina is a CSF that contains molecular interaction terms consisting of a van der Waals-like potential (defined by a combination of a repulsion term and two attractive Gaussians), a nondirectional hydrogen-bond term, a hydrophobic term, and a conformational entropy penalty, all of which are weighted by empirical parameters. The other three methods belong to MLSF category but vastly differ in their feature set. The RF-Score model is based on random forest regression that predicts binding affinities from the number of occurrences of a particular protein-ligand atom type pair interacting within a certain distance range using 3D structures. InteractionGraphNet (IGN) utilizes a graph neural network operating on the 3D complex structures, and the node and edge features are straightforward information about atoms and bonds, including atom types, atom hybridization, bond order, etc. Finally, DeepDTA is a Y-shaped 1-dimensional convolutional neural network that takes in protein sequences and ligand isomeric SMILES strings as input and outputs the predicted binding affinities, making it a model that does not rely on the exact 3D structure of the complex.
Details on the retraining of these models are described in the Supplementary Information. In short, AutoDock vina was retrained by optimizing its six empirical parameters by minimizing the mean absolute error between predicted and experimental binding affinities using the Nelder-Mead optimization algorithm81. Both IGN and DeepDTA were retrained by minimizing the mean squared error using gradient descent method as implemented in PyTorch82. For RF-Score, the optimization was driven by variance reduction between prediction and experimental affinities in each decision tree using the scikit-learn implementation83. After retraining, the new AutoDock Vina, RF-Score, IGN, and DeepDTA model performances are compared with the old models as tested on the non-covalent LP-PDBBind test set using the CL2 data, the BDB2020+ new benchmark data, the Mpro and EFGR applications, and finally evaluated on the redocked structures in BDB2020+ dataset.
@@ 3.1 Analysis of PDBBind Splittings
One of the main purpose of defining the new split of PDBBind is to prevent data leakage between training, validation, and test data. To understand whether the new split has solved the data leakage issue, the maximum similarity for proteins and ligands between the training, validation, and test data under the EquiBind split and LP-PDBBind, and maximum similarity between the general set and core set under the original PDBBind split are summarized in Figure 2(d-g). The original PDBBind split has significant protein and ligand overlap between the general and core sets, as can be seen in the sharp peak at similarity of 1.0 in Figure 2(d,f). The same level of similarity between the refined and core sets in the original PDBBind split are shown in Supplementary Figure 2. Since many ML models use the PDBBind core set as the test dataset without carefully excluding similar data from their respective training dataset, the performance of these models reported may be overly optimistic and do not reflect their true generalizability and needs to be reevaluated.
The Equibind split is a significant step forward in reorganizing the PDBBind data more reasonably. Its time-based cutoff in defining the test dataset decreased the chance of data
@@
leakage, but still did not eliminate the possibility of highly-similar data that occurs in both train and test set. By comparison, LP-PDBBind minimizes overlap between data in the train and test set by design, keeping the maximum sequence similarty between any protein in the test set and any protein in the train set below 0.5, and the maximum ligand similarity below 0.99 (Figure 2(d,f)). Therefore, results for a model trained with the LP-PDBBind train set and evaluated with LP-PDBBind test set should better reflect the performance of the model when applied to a new protein-ligand complex that may be very different than data used for training the model.
@@
Given that proteins with low sequence similarity might still be homologous and contain binding pockets that share similar interaction characteristics27, it is essential to double check on the similarities of protein-ligand interaction patterns for complexes in this new split of PDBBind dataset. The proteo-chemometrics based interaction fingerprints 49 were used to evaluate the interaction pattern similarities. Figure 2h shows the distribution of the maximum interaction fingerprint similarities for each complex in the test dataset with all complex structures in the training dataset under both the new LP-PDBBind split and the time-based Equibind split, and also the comparison between data in PDBBind core set and general set. LP-PDBBind has also eliminated data with high fingerprint similarity between train and test, which means the data with low protein sequence or ligand similarity but high interaction similarity is not an issue for the LP-PDBBind dataset, and the performance of any scoring function trained with LP-PDBBind train set and evaluated with the test set should be a realistic reflection of the model performance even when the protein-ligand interaction patterns are quite different from the complexes used for training.
@@
To help with training more transferable models, we have also defined the validation set to be equally different than the train set as the test set. Figure 2(e,g,i) illustrates the maximum protein similarities, maximum ligand similarities and maximum interaction fingerprint similarities between the validation and training sets for the Equibind split and LP-PDBBind, respectively. The similarity between validation and training sets under LP-PDBBind is also well controlled so that the validation set can be used to select the most transferable model or hyperparameters when training the model. By comparison, the validation set in Equibind split is too similar to the train set, hence overfitting will not be effectively captured when monitoring model performance on the validation set. As is provided in Supplementary Figure 3, both the Equibind and LP-PDBBind splits of PDBBind, the validation and test set have a wide range of (dis)/similarities on proteins, ligands and their interactions. Therefore using the validation set for model selection will not lead to overfitting and thus increase transferability to the test set.
@@ 3.2 Evaluating Retrained Models with LP-PDBBind
After the AutoDock Vina, IGN, RF-Score and DeepDTA models were retrained using LP-PDBBind, the performance were first compared on the LP-PDBBind test dataset filtered to CL2 for better reliability of data. Table 1 provides the mean and standard deviations on the root mean square error (RMSE) of the binding affinity prediction ($\Delta G_{bind}$) on the training, validation, and test data for the models retrained with LP-PDBBind from 3 different model initializations, and comparing it to the original models and their performance on the LP-PDBBind test data. Supplementary Figure 4 shows the scatter plots of the best model and reports the correlation coefficient between predicted and experimental binding affinities. Due to the data leakage issue, performance of the original models on the LP-PDBBind test data are over-estimated for the MLSFs. By comparison, AutoDock Vina due to its small number of trainable parameters does not suffer from the data leakage issue, and has achieved lower RMSE after retraining. Among the MLSF models, IGN has the smallest generalization gap, and is also the best performing model when evaluated using the LP-PDBBind test set within uncertainties. Since random forest models can almost perfectly fit training data 84, 85, the RF-Score model has exceptional performance on the LP-PDBBind training dataset, but its validation and test performance is on par with the IGN model. DeepDTA also performs quite well on the training dataset, but exhibits a large generalization gap with respect to the validation dataset, and has slightly higher RMSE on the test dataset than the other two MLSFs. Overall the original MLSF models have seen some of the LP-PDBBind test proteins and ligands in their training, and thus they appear to perform better than they actually do when data leakage is controlled for using the LP-PDBBind data.
Table 1: The training and test errors for the original and retrained models using LP-PDBBind for AutoDock Vina, IGN, RF-Score, and DeepDTA in terms of root mean square error (RMSE) in kcal/mol. Reported standard deviations were calculated with 3 different random initializations of model weights during training.
@@
DeepDTA 1.34 1.41$\pm$ 0.11 2.07$\pm$ 0.02 2.29$\pm$ 0.04
Figure 3 and Table 2 summarize the scoring performance of the original and retrained models on the independent BDB2020+ benchmark dataset; the corresponding scatter plots for the original and retrained models are provided in Supplementary Figure 5. We see that upon retraining, all four models have achieved meaningful improvements on the BDB2020+ benchmark set. In terms of RMSE, AutoDock Vina decreased by 1.2 kcal/mol while the other three MLSFs decreased by 0.2-0.3 kcal/mol. The changes in the correlation coefficients are more profound: both IGN and RF-Score have achieved absolute correlation coefficients better than 0.5, and improvements for AutoDock Vina and DeepDTA are also clear. Overall
Table 2: Performance comparisons on the BDB2020+ benchmark set in terms of root mean square error (RMSE) and Pearson correlation coefficient ($R$) for different models before and after retraining using LP-PDBBind. Reported standard deviations were calculated with 3 different random initializations of model weights during training.
@@
AutoDock Vina 3.31 2.10 $\pm$ 0 -37% 0.23 0.29$\pm$0
IGN 1.62 1.38$\pm$ 0.09 -9$\sim$200.41 0.54$\pm$0.04
RF-Score 1.80 1.61 $\pm$ 0.008 -10$\sim$110.36 0.51$\pm$ 0.008
DeepDTA 1.98 1.72$\pm$0.10 -9$\sim$190.18 0.26$\pm$0.07
@@
Overall, IGN and RFScore have improved more significantly than the other two, especially IGN that has achieved 1.38 $\pm$ 0.09 kcal/mol in RMSE and 0.54$\pm$ 0.04 in correlation coefficient when evaluated on the BDB2020+ dataset. We attribute the improvements on the model performances to our resplit of the PDBBind dataset because the training dataset becomes more diverse and representative, and that helps the SFs to find more transferable features and make more accurate predictions.
Figure 3 also compares the performance of the four models on BDB2020+ with the results evaluated on the LP-PDBBind test dataset. Interestingly, we find that all models have achieved lower RMSE on the BDB2020+ dataset but also lower Pearson correlation coefficients, $R$, other than IGN that has similar correlation coefficients between LP-PDBBind test dataset and BDB2020+ benchmark set. This seemingly contradictory result is actually reasonable due to the dataset distribution differences of the two evaluation benchmarks. The LP-PDBBind test dataset contains much more data than BDB2020+, and also spans a wider range of binding affinity values. The measured $-\log(K_{d})$ values in the LP-PDBBind test dataset ranges from 0 to 12 (i.e. 12 orders of magnitude), but the BDB2020+ dataset only ranges from 4 to 10. Given that extreme predictions from a robust ML model is unlikely, a narrower range of binding affinities means the overall error will be smaller. However, it also poses challenge for successfully differentiating the nuances between more clustered data points, and therefore it is also more difficult to achieve higher correlation coefficients.
@@ 3.3 Evaluating the Ranking Capabilities of the Retrained Models
Figure 4 shows the distributions of experimental binding affinities and protein and ligand similarities when comparing the LP-PDBBind training data with the Mpro and EGFR datasets. The binding affinities for these two systems are found to be in a narrower range than the LP-PDBBind training data. While the average binding affinity of the Mpro dataset is roughly in line with the LP-PDBBind training dataset, the ligands in the EGFR dataset have an overall tendency to bind stronger to their targets. This shift in the average indicates that the EGFR is more ”out-of-distribution” than the Mpro dataset. By calculating the Kullback-Leibler divergence (KL-divergence)86 with the kernel density estimations between binding affinities in the LP-PDBBind train dataset and the two target-specific datasets, we found that the KL-divergence between LP-PDBBind train dataset and SARS-COV-2 Mpro dataset is 3.45, and the KL-divergence between LP-PDBBind train dataset and EGFR dataset is 4.46, also validating the EFGR dataset has more dissimilar binding affinities to the training dataset. Figure 4b-e further show that the protein and sequence similarities of the two evaluation datasets are overall very dissimilar to the LP-PDBBind training dataset, with the exception of a small fraction of protein and ligand sequence similarity attributable to the SARS-CoV-1 Mpro training entry PDB ID: 3V3M87 for the Mpro dataset. Hence the Mpro and EFGR evaluations reflect two representative scenarios of using the SFs on a
@@
Figure 4: Data statistics for the SARS-CoV-2 main protease (Mpro) benchmark set and epidermal growth factor receptor (EGFR) benchmark set. (a) Distributions of the binding affinity data ($-\log K_{d}$) in the LP-PDBBind train dataset in blue, Mpro benchmark set in orange and EGFR set in green. (b-c) Distributions of protein sequence similarities between the Mpro protein (b) and EGFR protein (c) with proteins in the LP-PDBBind train dataset. (d-e) Distributions of ligand fingerprint similarities between molecules in the Mpro benchmark set (d) and EGFR benchmark set (e) with ligands in the LP-PDBBind train dataset.
similar or different protein-ligand system than what wasincluded in the PDBBind dataset.
@@
Table 3 summarizes the RMSE, Pearson correlation coefficients ($R$) and Spearmann correlation coefficients ($R_{S}$ or ranking power) for the data in the Mpro and EGFR evaluation sets using both the original models and the models retrained using LP-PDBBind. The prediction scatter plots for the Mpro benchmark dataset and EGFR benchmark dataset are provided in Supplementary Figures 6 and 7. For the Mpro dataset, all four models exhibit some modest improvement in terms of RMSE. AutoDock Vina, RF-Score and DeepDTA have all achieved higher correlations in terms of $R$ and $R_{S}$. The correlation on the retrained IGN model has slightly decreased but is still similar to the original values. Taken all metrics into account, AutoDock Vina is the best performing SF for the Mpro dataset, although the differentiation among all models is not large.
@@
However, the results on the EGFR dataset exhibit much larger variations for the different models, and the changes are significant for all the SFs. There is a dramatic decrease by ~50% in RMSE for AutoDock Vina, and both $R$ and $R_{S}$ have increased significantly. IGN has a slight decrease in RMSE but the correlation coefficients have improved remarkably from 0.36 to 0.65 in Pearson correlation coefficient and from 0.17 to 0.62 in in Spearmann correlation coefficient. These values are the highest among all retrained models, and do not differ much with the Mpro results and thus showing good generalizability. RF-Score has also benefitted from retraining quite significantly, with a 0.6 kcal/mol decrease in RMSE, and the correlation coefficients have improved from -0.15 to 0.52 in terms of $R$ and from -0.18 to 0.45 in terms of $R_{S}$. Finally, DeepDTA has also seen improvements upon retraining, but the ranking capability of the retrained DeepDTA model is still inferior to the other two MLSFs. Overall, the performance rankings of the four SFs are still consistent with that obtained from the LP-PDBBind test set benchmark and the BDB 2020+ benchmark. But the EFGR benchmark emphasizes that new applications will benefit from better generalizability of the newly retrained models, especially for IGN.

### v3

@@ 2.5 Preparation of two target-specific datasets
Finally, we have prepared two additional datasets of protein-ligand complexes with the same protein and different ligands so that we could evaluate the ranking accuracies of the SFs before and after retraining using LP-PDBBind. The first dataset is based on the SARS-CoV-2 main protease (Mpro) for which a wide variety of potential Mpro inhibitors have been developed.13 We have manually extracted published co-crystal structures of Mpro with a number of non-covalent inhibitors68, 41, 66, 20, 10, 30, 83, 16, 8, 26, 31, 14, 23, and prepared a dataset containing 40 structures and corresponding experimental binding affinity measurements. The second dataset involves the epidermal growth factor receptor (EGFR), which is a receptor tyrosine kinase related to multiple cancers including lung cancer, pancreatic cancer and breast cancer.25 Similarly, we have selected 23 representative non-covalent protein-ligand complex structures of EGFR with binding affinities taken from BindingDB.80, 85, 76, 53, 24, 62, 64, 5, 54, 1, 56, 81, 75, 78, 74 The two dataset were selected as two representative scenarios of using the SFs on a similar or different protein-ligand complex than what was included in the training dataset.
@@ 2.6 Retraining Scoring Functions with LP-PDBBind
The LP-PDBBind data for non-covalent binders of PDBBind were used to retrain AutoDock Vina67, IGN28, the 2010 RF-Score2 and DeepDTA50. These models cover a wide range of different approaches for scoring, and are representative of different dimensions of the SF space of models. AutoDock vina is a CSF that contains molecular interaction terms consisting of a van der Waals-like potential (defined by a combination of a repulsion term and two attractive Gaussians), a nondirectional hydrogen-bond term, a hydrophobic term, and a conformational entropy penalty, all of which are weighted by empirical parameters. The other three methods belong to MLSF category but vastly differ in their feature set. The RF-Score model is based on random forest regression that predicts binding affinities from the number of occurrences of a particular protein-ligand atom type pair interacting within a certain distance range using 3D structures. InteractionGraphNet (IGN) utilizes a graph neural network operating on the 3D complex structures, and the node and edge features are straightforward information about atoms and bonds, including atom types, atom hybridization, bond order, etc. Finally, DeepDTA is a Y-shaped 1-dimensional convolutional neural network that takes in protein sequences and ligand isomeric SMILES strings as input and outputs the predicted binding affinities, making it a model that does not rely on the exact 3D structure of the complex.
Details on the retraining of these models are described in the Supplementary Information. In short, AutoDock vina was retrained by optimizing its six empirical parameters by minimizing the mean absolute error between predicted and experimental binding affinities using the Nelder-Mead optimization algorithm48. Both IGN and DeepDTA were retrained by minimizing the mean squared error using gradient descent method as implemented in PyTorch51. For RF-Score, the optimization was driven by variance reduction between prediction and experimental affinities in each decision tree using the scikit-learn implementation52. After retraining, the new AutoDock Vina, RF-Score, IGN, and DeepDTA model performances are compared with the old models as tested on the non-covalent LP-PDBBind test set using the CL2 data, the BDB2020+ new benchmark data, the Mpro and EFGR applications, and finally evaluated on the redocked structures in BDB2020+ dataset.
@@ 3.1 Analysis of PDBBind Splittings
One of the main purpose of defining the new split of PDBBind is to prevent data leakage between training, validation, and test data. To understand whether the new split has solved the data leakage issue, the maximum similarity for proteins and ligands between the training
@@
validation, and test data under the EquiBind split and LP-PDBBind, and maximum similarity between the general set and core set under the original PDBBind split are summarized in Figure 1(d-g). The original PDBBind split has significant protein and ligand overlap between the general and core sets, as can be seen in the sharp peak at similarity of 1.0 in Figure 1(d,f). The same level of similarity between the refined and core sets in the original PDBBind split are shown in Supplementary Figure S2. Since many ML models use the PDBBind core set as the test dataset without carefully excluding similar data from their respective training dataset, the performance of these models reported may be overly optimistic and do not reflect their true generalizability and needs to be reevaluated.
@@
The Equibind split is a significant step forward in reorganizing the PDBBind data more reasonably. Its time-based cutoff in defining the test dataset decreased the chance of data leakage, but still did not eliminate the possibility of highly-similar data that occurs in both train and test set. By comparison, LP-PDBBind minimizes overlap between data in the train and test set by design, keeping the maximum sequence similarty between any protein in the test set and any protein in the train set below 0.5, and the maximum ligand similarity below 0.99 (Figure 1(d,f)). Therefore, results for a model trained with the LP-PDBBind train set and evaluated with LP-PDBBind test set should better reflect the performance of the model when applied to a new protein-ligand complex that may be very different than data used for training the model.
@@
Given that proteins with low sequence similarity might still be homologous and contain binding pockets that share similar interaction characteristics84, it is essential to double check on the similarities of protein-ligand interaction patterns for complexes in this new split of PDBBind dataset. The proteo-chemometrics based interaction fingerprints 70 were used to evaluate the interaction pattern similarities. Figure 1h shows the distribution of the maximum interaction fingerprint similarities for each complex in the test dataset with all complex structures in the training dataset under both the new LP-PDBBind split and the time-based Equibind split, and also the comparison between data in PDBBind core set and general set. LP-PDBBind has also eliminated data with high fingerprint similarity between train and test, which means the data with low protein sequence or ligand similarity but high interaction similarity is not an issue for the LP-PDBBind dataset, and the performance of any scoring function trained with LP-PDBBind train set and evaluated with the test set should be a realistic reflection of the model performance even when the protein-ligand interaction patterns are quite different from the complexes used for training.
@@
To help with training more transferable models, we have also defined the validation set to be equally different than the train set as the test set. Figure 1(e,g,i) illustrates the maximum protein similarities, maximum ligand similarities and maximum interaction fingerprint similarities between the validation and training sets for the Equibind split and LP-PDBBind, respectively. The similarity between validation and training sets under LP-PDBBind is also well controlled so that the validation set can be used to select the most transferable model or hyperparameters when training the model. By comparison, the validation set in Equibind split is too similar to the train set, hence overfitting will not be effectively captured when monitoring model performance on the validation set. As is provided in Supplementary Figure S3, both the Equibind and LP-PDBBind splits of PDBBind, the validation and test set have a wide range of (dis)/similarities on proteins, ligands and their interactions. Therefore using the validation set for model selection will not lead to overfitting and thus increase transferability to the test set.
@@ 3.2 Evaluating Retrained Models with LP-PDBBind
After the AutoDock Vina, IGN, RF-Score and DeepDTA models were retrained using LP-PDBBind, the performance were first compared on the LP-PDBBind test dataset filtered to CL2 for better reliability of data. Table 1 provides the mean and standard deviations on the root mean square error (RMSE) of the binding affinity prediction ($\Delta G_{bind}$) on the training, validation, and test data for the models retrained with LP-PDBBind from 3 different model initializations, and comparing it to the original models and their performance on the LP-PDBBind test data. Supplementary Figure S4 shows the scatter plots of the best model and reports the correlation coefficient between predicted and experimental binding affinities.
Table 1: Training and test errors for the original and retrained models using LP-PDBBind. Results for AutoDock Vina, IGN, RF-Score, and DeepDTA in terms of root mean square error (RMSE) in kcal/mol. Reported standard deviations were calculated with 3 different random initializations of model weights during training.
@@
[…]
Due to the data leakage issue, performance of the original models on the LP-PDBBind test data are over-estimated for the MLSFs. By comparison, AutoDock Vina due to its small number of trainable parameters does not suffer from the data leakage issue, and has achieved lower RMSE after retraining. Among the MLSF models, IGN has the smallest generalization gap, and is also the best performing model when evaluated using the LP-PDBBind test set within uncertainties. Since random forest models can almost perfectly fit training data 7, 3, the RF-Score model has exceptional performance on the LP-PDBBind training dataset, but its validation and test performance is on par with the IGN model. DeepDTA also performs quite well on the training dataset, but exhibits a large generalization gap with respect to the validation dataset, and has slightly higher RMSE on the test dataset than the other two MLSFs. Overall the original MLSF models have seen some of the LP-PDBBind test proteins and ligands in their training, and thus they appear to perform better than they actually do when data leakage is controlled for using the LP-PDBBind data.
[…]
Table 2: Performance comparisons on the BDB2020+ benchmark set. Root mean square error (RMSE) and Pearson correlation coefficient ($R$) for different models before and after retraining using LP-PDBBind. Reported standard deviations were calculated with 3 different random initializations of model weights during training.
@@
[…]
Figure 2 also compares the performance of the four models on BDB2020+ with the results evaluated on the LP-PDBBind test dataset. Interestingly, we find that all models have achieved lower RMSE on the BDB2020+ dataset but also lower Pearson correlation coefficients, $R$, other than IGN that has similar correlation coefficients between LP-PDBBind test dataset and BDB2020+ benchmark set. This seemingly contradictory result is actually reasonable due to the dataset distribution differences of the two evaluation benchmarks. The LP-PDBBind test dataset contains much more data than BDB2020+, and also spans a wider range of binding affinity values. The measured $\log(K_{d})$ values in the LP-PDBBind test dataset ranges from 0 to 12 (i.e. 12 orders of magnitude), but the BDB2020+ dataset only
@@
ranges from 4 to 10. Given that extreme predictions from a robust ML model is unlikely, a narrower range of binding affinities means the overall error will be smaller. However, it also poses challenge for successfully differentiating the nuances between more clustered data points, and therefore it is also more difficult to achieve higher correlation coefficients.
@@ 3.3 Evaluating the Ranking Capabilities of the Retrained Models
Figure 3 shows the distributions of experimental binding affinities and protein and ligand similarities when comparing the LP-PDBBind training data with the Mpro and EGFR datasets. The binding affinities for these two systems are found to be in a narrower range than the LP-PDBBind training data. While the average binding affinity of the Mpro dataset is roughly in line with the LP-PDBBind training dataset, the ligands in the EGFR dataset
@@
Figure 3: Data statistics for the SARS-CoV-2 main protease (Mpro) benchmark set and epidermal growth factor receptor (EGFR) benchmark set. (a) Distributions of the binding affinity data ($\log K_{d}$) in the LP-PDBBind train dataset in blue, Mpro benchmark set in orange and EGFR set in green. (b-c) Distributions of protein sequence similarities between the Mpro protein (b) and EGFR protein (c) with proteins in the LP-PDBBind train dataset. (d-e) Distributions of ligand fingerprint similarities between molecules in the Mpro benchmark set (d) and EGFR benchmark set (e) with ligands in the LP-PDBBind train dataset.
have an overall tendency to bind stronger to their targets. This shift in the average indicates that the EGFR is more ”out-of-distribution” than the Mpro dataset. By calculating the Kullback-Leibler divergence (KL-divergence)32 with the kernel density estimations between binding affinities in the LP-PDBBind train dataset and the two target-specific datasets, we found that the KL-divergence between LP-PDBBind train dataset and SARS-COV-2 Mpro dataset is 3.45, and the KL-divergence between LP-PDBBind train dataset and EGFR dataset is 4.46, also validating the EFGR dataset has more dissimilar binding affinities to the training dataset. Figure 3b-e further show that the protein and sequence similarities of the two evaluation datasets are overall very dissimilar to the LP-PDBBind training dataset, with the exception of a small fraction of protein and ligand sequence similarity attributable to the SARS-CoV-1 Mpro training entry PDB ID: 3V3M27 for the Mpro dataset. Hence the Mpro and EFGR evaluations reflect two representative scenarios of using the SFs on a similar or different protein-ligand system than what wasincluded in the PDBBind dataset.
@@
Table 3 summarizes the RMSE, Pearson correlation coefficients ($R$) and Spearmann correlation coefficients ($R_{S}$ or ranking power) for the data in the Mpro and EGFR evaluation sets using both the original models and the models retrained using LP-PDBBind. The prediction scatter plots for the Mpro benchmark dataset and EGFR benchmark dataset are provided in Supplementary Figures S6 and S7. For the Mpro dataset, all four models exhibit some modest improvement in terms of RMSE. AutoDock Vina, RF-Score and DeepDTA have all achieved higher correlations in terms of $R$ and $R_{S}$. The correlation on the retrained IGN model has slightly decreased but is still similar to the original values. Taken all metrics into account, AutoDock Vina is the best performing SF for the Mpro dataset, although the differentiation among all models is not large.
@@
However, the results on the EGFR dataset exhibit much larger variations for the different models, and the changes are significant for all the SFs. There is a dramatic decrease by ˜50% in RMSE for AutoDock Vina, and both $R$ and $R_{S}$ have increased significantly. IGN has a slight decrease in RMSE but the correlation coefficients have improved remarkably from 0.36 to 0.65 in Pearson correlation coefficient and from 0.17 to 0.62 in in Spearmann correlation coefficient. These values are the highest among all retrained models, and do not differ much with the Mpro results and thus showing good generalizability. RF-Score has also benefitted from retraining quite significantly, with a 0.6 kcal/mol decrease in RMSE, and the correlation coefficients have improved from -0.15 to 0.52 in terms of $R$ and from -0.18 to 0.45 in terms of $R_{S}$. Finally, DeepDTA has also seen improvements upon retraining, but the ranking capability of the retrained DeepDTA model is still inferior to the other two MLSFs. Overall, the performance rankings of the four SFs are still consistent with that obtained from the LP-PDBBind test set benchmark and the BDB 2020+ benchmark. But the EFGR benchmark emphasizes that new applications will benefit from better generalizability of the newly retrained models, especially for IGN.